OpenStudy (anonymous):

the number of cubic feet in the volume of a cube is the same as the number of square feet in its surface area. the length of the edge of the cube, in feet is?

OpenStudy (anonymous):

Set up the volume and surface area formulas in terms of side length and set the lengths equal.

OpenStudy (anonymous):

V, Volume = s^3 A, Surface area = 6s^2 V=A --> s^3 = 6s^2.